20230223_SX_DX_Night DX S-LUCI Comm. Closed: bad weather.

Startup System

07:00 Power on DX WFS

632 modes

7:44 DX AdSec in Failure.
  • Had just set it, flat for ~30 seconds
  • AdSecArbitrator_62424|INF| 228|2023-02-24 07:44:11.635537| COMMANDHANDLER > Command Set Flat AO (code 2009) successfully completed
  • AdSecArbitrator_62424|ERR| 229|2023-02-24 07:44:41.638846| ALERTHANDLER > Received an ERROR alert! Forcing Failure state.
  • AdSecArbitrator_62424|ERR| 230|2023-02-24 07:44:41.638904| ALERTHANDLER > Alert message: COILS DISABLED from housekeeper.R
  • adsc_stop/start

7:52 DX AdSec Set again after stop/start
  • 632 modes
  • Forces -0.28 to 0.22
  • No Problems

7:55 Preset with 10 mag
  • bin 1, 1500 Hz, 10.9 mag
  • 1.0" seeing
  • 080725, 080755, 080822

8:19 LUCI interface up and running

8:21 Link to new AO Table
  • [AOeng@soul-dxwfs ao]$ ln -sib table_LUCIFER_ACE-AO_20230223_500m_v02.txt table_LUCIFER_ACE-AO.txt
    • ln: replace `table_LUCIFER_ACE-AO.txt'? yes
  • Feb 24 08:20 table_LUCIFER_ACE-AO.txt -> table_LUCIFER_ACE-AO_20230223_500m_v02.txt

8:23 Preset bin 4
  • 16.0 mag
  • light source at 95. will adjust
  • 1.5" seeing @ 200 Hz
  • 277 Hz
  • LUCI J+Clear, 20 sec, 3 frames
  • 084036 (full LUCI Frames)
  • 084408, 084513, 084607, 084645, 084726, 084809 (512x512 subwin)
  • Dark 084949, 085105

9:13 Fabio add OCAM routines to WFS arbitrator

10:06 If zero slopes then run bcu dump script then the reset_gopt script
  • /home/AOeng/live/source/scripts/bcuDumpScript.py
  • /home/AOeng/live/source/reset_gopt.sh

10:06 If negative OCAM intensity
  • /home/AOeng/live/source/scripts/bcuDumpScript.py

10:35 Fabio will make a few additions and then make request to LBTO deploy new AUXLOOP

10:35 Doug will send link to document and instructions to LBTO Users to take bcu Dumps and send email to SOUL Team

10:47 Preset 17.3 mag
  • bin 4, 200 Hz, 1.5" seeing @ 200 Hz
  • gain 0.25, 1.31
  • Enrico change Optimize gain parameters
  • gain 2.97, 1.29
  • GOPT: 0.16
  • TN 110809, 110925, 111000 (short exp time, forgot to reset after stop/start)
  • TN 111211 (3 frames @ 20 sec)
  • TN 111353, 111602, 111813, 112025, 112229 (5 frames @ 20 sec)
  • Dark 112703, 112916

11:35 Preset 18.2 mag
  • bin 4, 134 Hz, 18.1 mag
  • gain 1.59, 0.56
  • GOPT 0.35
  • TN 114054, 114318, 114524, 114731, 114933

11:55 Done for tonigth

-- DougMiller - 24 Feb 2023
Topic revision: r2 - 24 Feb 2023, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback